Solve the system
x-1y= 3
3x+2y= 14

Answer by arunpaul(104)

x-1y= 3
3x+2y= 14


----------
3x-3y=9
3x+2y=14
-----------
subtracting we have
-5y=-5
so y =1
so x = 4 by putting the value in any equation
(Ans)